Adding unity to the numerator as well as the denominator of a fraction makes it equal to 4/5. Subtracting 5 from each makes it equal to 1/2. What is the fraction?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Adding unity to the numerator as well as the denominator of a fraction makes it equal to 4/5. Subtracting 5 from each makes it equal to 1/2. What is the fraction?
---------------
Equations::
(x+1)/(y+1) = 4/5
(x-1)/(y-1) = 1/2
----------
Cross-multiply to get:
5x+5 = 4y+4
2x-2 = y-1
-------
Rearrange::
5x - 4y = -1
2x - y = 1
------
Modify for elimination::
5x - 4y = -1
8x - 4y = 4
-----------------
Subtract and solve for "x"::
3x = 5
x = 5/3
----
Solve for "y"
2x-y = 1
10/3 - y = 1
y = 7/3
------
Ans: # = (5/3)/(7/3) = 5/7
-------------------------------
Cheers,
Stan H.
-------------

Answer by richwmiller(17219)

(x+1)/(y+1) = 4/5,
(x-5)/(y-5) = 1/2
5*(x+1)=4*(y+1)
5x+5=4y+4
(x-5)/(y-5) = 1/2
2*(x-5)=(y-5)
2x-10=y-5
2x=y+5
y=2x-5
5x+5=4y+4
5x+5=4(2x-5)+4
5x+5=8x-20+4
5x+5=8x-16
3x=21
x = 7
y=2x-5
y-14-5
y = 9
7/9 is original fraction